OpenStudy (anonymous):

A bug is hovering in the air when a vehicle moving at 20m/s hits it. The colision is perfectly elastic, what is the speed of the bug after collision

OpenStudy (anonymous):

Perfectly elastic collision = No change in kinetic energy (conservation of Kinetic energy?)

OpenStudy (anonymous):

\[KE_i=KE_f\]

OpenStudy (anonymous):

for an elastic collision with a stationary target \[m _{1} V _{1i}^{} = m _{1} *V _{1f}^{} +m _{2} *V _{2f}^{} \] use: \[V _{2f}^{} = \frac{ 2m _{1} }{ m _{1} +m _{2}}*V _{1i}^{} \]

OpenStudy (anonymous):

\[\frac{1}{2}m_cv_c^2=\frac{1}{2}m_bv_b'^2+\frac{1}{2}m_cv_c'^2 \] \[m_cv_c=m_bv_b'+m_cv_c'\]

OpenStudy (anonymous):

\[\frac{m_cv_c-m_cv'_c}{m_b}=v'_b \] \[\frac{2m_c}{m_c+m_b}*v_b=v_b'\]

OpenStudy (anonymous):

@Algebraic! whered you come up with 2m at the top

OpenStudy (anonymous):

You equation itself can't be correct because i'm not given any masses so i can't solve it... somewhere mass must cancel

OpenStudy (anonymous):

k.

OpenStudy (anonymous):

m2~0 now try it

OpenStudy (anonymous):

going to need two simultaneous equations 1) Kinetic energies will equal the same which i did now

OpenStudy (anonymous):

that makes sense but how'd you come to that conclusion

OpenStudy (anonymous):

it comes from using KEi =KEf to simplify m1v1i = m1v1f +m2v2f

OpenStudy (anonymous):

also you give no explaination of how you came to that equation.... .was it from simultaneously solving both the kinetic and the momentum equations?

OpenStudy (anonymous):

derivation is probably shown in your text. it's just algebra. I've already done it for people several times on this site and it's invariably a waste of time, because they always say "oh, that was in my book actually" after you've gone through the steps for them.